OpenStudy (anonymous):

PLEASE HELP MEDAL WILL BE GIVEN NEED HELP ASAP. write the sum using summation notation, assuming the suggested pattern continues -1+2+5+8+...+44

OpenStudy (anonymous):

@aaronq @Compassionate @confluxepic

OpenStudy (anonymous):

@ganeshie8 @mathstudent55

OpenStudy (kirbykirby):

Hint: \(-1+3(0)=-1 \\ -1+3(1)=2\\ -1+3(2)=5\\ -1+3(3)=8 ... \) Do you see the pattern? Notice how you can replace the number that keeps changing by a variable to be used in summation notation

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}(-1+3n)\]

OpenStudy (kirbykirby):

That is almost it! :) The sum ends at 44 though, it doesn't continue till infinity, so you need to figure out the number at which is stops. This can easily be done with a small equation: \(-1+3n=44 \)

OpenStudy (anonymous):

\[\sum_{n=0}^{15}(-1+3n)\]

OpenStudy (kirbykirby):

awesome :)!!

OpenStudy (anonymous):

sweet okay thank you

OpenStudy (kirbykirby):

np